OpenStudy (anonymous):

-t/16=-7/8 what is t, please show work.

OpenStudy (anonymous):

-t = (16 *(-7))/8 t = 14

OpenStudy (anonymous):

clear enough ?

OpenStudy (anonymous):

well, were supposed to keep what is on each side, on each side and solve like that.

OpenStudy (anonymous):

well you can multiply the right side by 0.5/0.5 which is 1 so everything is fine and then this side will become : -0.5t/8 now you will see -0.5t = -7 t = 14 which is the same as before but maybe this is what you need

OpenStudy (amistre64):

in other words, dont solve for t. use a proscribed method and determine t by that method.

OpenStudy (amistre64):

the question should be: prove that t = 14 by using a few different methods.

OpenStudy (anonymous):

t=14